The CDF Experiment

Study of p p interactions at √s = 1.8 TeV

• 22 institutions

• 230 physicists

IPP Group consists of

McGill University  University of Toronto
faculty  1  1
research associate  1  1
graduate students  2  2

Completed analysis of 1988-89 data set

• Preparing for run in 1992-93 (Run IA and IB)

• Expect to record 100 pb−1

Top Quark Search

Electron + Nuetrino + Jets Analysis

• electron + neutrino ID eliminates QCD background • Remaining background from W + jets production

• W + jets subtracted statistically using transverse mass

⇒ 40 < Mtop < 77 GeV/c2 (95% CL)

Electron + Muons

• e±μ∓ is the cleanest channel for top search

• observed only 1 event in data sample

⇒ 28 < Mtop < 72 GeV/c2 (95% CL)

Recent top mass limit comes from combining channels

• Lepton + Jets – Required b → μX candidate in event

– Added μ + jets channel (×2)

– No events pass selection (t t efficiency is ∼ 0.2%)

• Dileptons – Added e+ e− and μ+ μ− channels

– Rejected Z∘ with M(l+ l−) cut

– Rejected Drell Yan using Δϕ and E̸T cut

– 70% increase in dilepton acceptance for t t

• Only one event survives the selection

⇒ 28 < Mtop < 91 GeV/c2 (95% CL)

Bottom Quark Physics

Bottom Quark Production Cross Section

• Measured σb using – b → e+ νeX

– B → J/ΨK±

• Inclusive e± sample dominated by b production – background from QCD < ∼ 15%

– charm quark background rejected by isolation cuts ∗ estimated to be 10 ± 10%

• B → J/ΨK± analysis based on signal of ∼ 15 events – Systematic uncertainties are quite different

⇒ σb is ×2 larger than theory predicts

B^{∘} − B̄^{∘} Mixing

• Mixing analysis based on 911 low-energy e − μ events

R = \frac{N(e+ μ+ ) + N(e−μ−)}{N(e+ μ−) + N(e−μ+ )}

• Analysis corrected for – misidentified leptons (mostly fake μ’s)

– c quark contribution

– sequential decays

• Expressed result in terms of the mixing parameter χ

χ ≡ Prob (B∘ → B̄∘)
= 0.176 ± 0.026(stat ) ± 0.025(bkg ) ± 0.032(MC)

Other B physics analyses underway:

1. B → J/ΨK∘ s

2. Upper limit on B → μ+ μ−

3. Search for Λb

4. B lifetime studies

CDF has been studying a number of other phenomena

1. W and Z∘ masses

2. Production cross sections and asymmetries for W and Z∘ 3. Search for additional bosons (W ′’s and Z′’s)

4. QCD tests

5. SUSY particle searches

Plans for 1992 CDF Run

Expect to obtain ∼ 100 pb^{−1}

• Run IA (6 months ⇒ 25 pb−1)

• Run IB (6 months ⇒ 75 pb−1)

• expect L∼ 7 × 1030 cm−2s−1

CDF Detector Upgrades

• Central muon upgrade and extension to |η| < 1.0

• SVX Detector (4 layers of 55 and 65 μ pitch Si strips)

• New inner vertex time projection chamber (VTX)

• Preradiator between coil and CEM (γ/π∘ ID)

• DAQ upgrade

• L3 SGI farm (⇒> 1200 MIPS)

• Offline and analysis software ported to UNIX

Physics Goals

Top Quark Search

• Lepton + jets mode requires b quark tagging – SVX

– b → μX and b → eX

• e±μ∓ channel expected to be quite clean

⇒ top quark mass “reach” of ∼ 150 − 160 GeV/c2

Measurement of M_{W}

• Hope to measure to ∼ 100 MeV/c2

• Limited by systematics (currently ∼ 250 MeV/c2)

Bottom Quark Physics

• Bs search and mixing

• b and τ lifetimes

• B rare decays

• Proof-of-principle of high-sensitivity collider B physics

CDF/IPP Activities in Last Year

DAQ Upgrade for 1992 run

• developed and commissioned new Event Builder

• modelled DAQ system and evaluated system performance • new system now in routine operation

• DAQ readout rate increased from 7 to ∼ 35 Hz

Offline Software Development

• simulation and analysis code ported to UNIX

• responsible for “public” software releases

• collaborated with FNAL on developing UNIX software distribution system

– have Toronto, McGill, Penn and FNAL code maintained “automatically”

Detector Simulation

• Ported package to UNIX

• implemented detector upgrades

• currently improving calorimeter simulation

Research for 1992-93

Group’s Activities will focus on:

• DAQ upgrade for Run IB

• Software development and management

• Detector simulation

• Data taking and physics analysis

DAQ Upgrade

• Require DAQ upgrade between Run IA and IB

• replace EVB and control system with VME-based event collection

• new DAQ interface to L3 and control system

• FE electronics will be replaced after Run IB

• IPP responsible for L3 interface and control system

• Currently prototyping processors, links and O/S

Software Development

• Continued development of detector simulation

• UNIX port of reconstruction appears to be under control • Focus shifting to analysis software

Physics Analysis Effort

• Resources to mount effective effort at Toronto and McGill • Will require consistent presence at FNAL

• Focus will be on top and bottom physics

Summary of Research Support

Group has 2nd year of 2-year operating grant for 92-93

• Provides barely minimum support for ongoing activities

• Not enough resources for DAQ upgrade

Requested Equipment Grant for the DAQ upgrade project

• Crucial to our effort

• No other group in CDF with requisite expertise and manpower

CDF DAQ Upgrade

Fermilab Collider is planning lengthy 1992-93 run

• 6 month run from 3/92—8/92 (25 pb−1)

• 6 month shutdown for Collider upgrades

• 6-8 month run from 3/93—9/93 (75 pb−1)

• Lmax ∼ 7 × 1030 → 15 × 1030 cm −2s−1

In 1988-89 run,

• Collider Lmax ≃ 2 × 1030

• CDF FE electronics and DAQ limited L2 trigger rate – < ∼ 5Hz for 5% FE deadtime – < ∼ 7Hz for 5% DAQ deadtime – < ∼ 5Hz L3 processing rate

⇒ σL2 < ∼ 2μb

For Run IA, have made a number of upgrades

• FE digitizition time reduced to < 2 ms

• added 2nd EVB and increased parallelization

• upgraded L3 farm from 50 → 1200 MIPS

⇒ maximum readout rate is > ∼ 30 Hz

Have not increased maximum L2 trigger cross section

• Makes addition of “b-specific” triggers difficult

• Leaves little margin if Lmax grows

In addition, entire FE and DAQ has to be replaced after Run IB

• Collider goes to 36-bunch operation

• Implement pipelined FE and L1 trigger

• DAQ system has to take 100 Hz L2 rate

SO...

A partial DAQ upgrade between Run IA and IB is being proposed

• Implement system with 50 Hz readout rate

• Provide additional margin for Run IB

• Easily scale to 100 Hz (and Run II)

CDF DAQ Upgrade

New system would consist of

• New Fastbus-based readout controller (FRC)

• Scanner CPU (SCPU) to interface FRC and L3

• UltraNet to ship data from SCPU to L3 (100 MB/s)

• VME-based control system for FRC, SCPU and L3

Philsophy of Upgrade

• Use as much commercial hardware as practical

• Dedicated data and control networks

• Avoid complex handshakes and control protocols

Responsibility for project would be shared

• FNAL responsible for design and implementation (including software) of FRC

• IPP responsible for design and implementation of SCPU • IPP responsible for VME-based control network

• FNAL and Rutgers responsible for UltraNet implementation • FRC-SCPU bus joint responsibility of IPP and FNAL

Budget for IPP share

Processors for Control and SCPU  26K
VxWorks Operating System Software  33K
Reflective memories  25K
Components for FRC-SCPU bus  21K
4 VME crates  11K
Cables, miscellaneous  10K
Total (approximate)  126K

Manpower

• P. Sinervo (0.25 FTE)

• K. Ragan (0.3 FTE)

• K. Biery (0.5 FTE)

• A. Holscher (0.5 FTE)

• R. Grindley (0.5 FTE)

Schedule set by Collider Run

• Prototyping of system currently underway – Evaluating VxWorks on 68030 board

– Acquired (with ZEUS) 2 reflective memories – FNAL evaluating UltraNet

– Verilog simulation of system almost complete • Expect to have complete design by 4/92

– FRC design review to take place 1/92

– VME control network design by 3/92

• Implementation to begin in 9/92 after Run IA